# Network Structure Trumps Neuron Dynamics: Insights from Drosophila Connectome Simulations

### v3

@@ 2.4 Distances between Input and Activation Areas
To further investigate how the network structure, rather than the spatial distribution of neurons, or the parameters of the network model or activation mechanisms, predominantly determines the dynamics of brain activity, we calculated the average network distance and spatial distance between different areas of the brain network. The optic and visual_projection area are related to vision and ALLN, ALIN, ALON, ALPN, LHCENT and LHLN are related to olfactory. The results are presented in Tab. 2.
From Tab. 2, it is evident that neurons which are closer in spatial distance are not necessarily easier to activate. For example, all areas and visual input neurons have similar average spatial distances, but the quasi-real activation patterns can only be observed in specific neural types (optic, visual_projection). In contrast, the average network distance between input neurons and neurons should be activated is closer than irrelevant neurons. For instance, the distance between visual input neurons and optic, visual_projection are 5.7018 and 5.7432, respectively. And the distance between visual input neurons and ALLN, ALIN, ALON, ALPN, LHCENT, LHLN which are irrelevant with vision are 6.6442, 6.2350, 6.1506, 6.5380, 6.3094 and 6.8206, respectively. The network distance of neurons related to visual input neurons is generally closer than the distances of unrelated neurons. Besides, The average network distance of neurons related to olfactory input neurons is approximately twice as close as the distance of unrelated neurons. These results indicates that compared to spatial distance, the average network distance between input neurons and their corresponding activation areas is closer, meaning that compared to spatial structure, network structure is more important, determining these quasi-real activation patterns of the brain network.
This results reflect the importance of network structure in neural networks, that is, the ability to reproduce certain functions of the brain by considering the network structure only.
Table 2: Spatial and network distance between input and main areas. The unit of spatial distance is nm and the unit of network distance is the number of edges in the shortest path.
distance type Neuron type Area Name optic visual_projection ALLN ALIN
Spatial Distance Visual Input Neurons $365012$ $354725$ $351919$ $348901$
Olfactory Input Neurons $302690$ $251144$ $110827$ $173345$
ALON ALPN LHCENT LHLN
Visual Input Neurons $346925$ $348517$ $362825$ $366013$
Olfactory Input Neurons $123179$ $115089$ $212515$ $246719$
optic visual_projection ALLN ALIN
Network Distance Visual Input Neurons $5.7018$ $5.7432$ $6.6442$ $6.2350$
Olfactory Input Neurons $5.7967$ $5.1706$ $3.4458$ $3.3845$
ALON ALPN LHCENT LHLN
Visual Input Neurons $6.1506$ $6.5380$ $6.3094$ $6.8206$
Olfactory Input Neurons $3.6312$ $3.6820$ $3.9557$ $3.9758$
@@ 3 Discussion
This study, through the construction of a large-scale network communication model based on the Drosophila connectome, reveals the critical role of neural network structure in generating brain-like activation patterns. Our findings suggest that models based on comprehensive neuron network data can approach the brain’s quasi-real activation patterns even with simple artificial rules. This paves the way for understanding biological intelligence, indicating that the structure of biological neural systems may offer valuable reference for designing efficient, fast-adapting intelligent systems.
Our research emphasizes the importance of network structure, compared to spatial structure or neuronal activation mechanisms, in determining brain activity patterns. This finding has significant implications for artificial neural network research, suggesting that future AI designs may need to more closely mimic real brain network structures.
Furthermore, the large spatial network visualization software we developed provides a new tool for neural network research, facilitating intuitive observation and analysis of complex neural network dynamics.
@@
Future research directions may include optimizing propagation rules and automatically adjusting link weights to achieve complete simulation of brain behavior. These advancements will pave the way for achieving genuine artificial intelligence while deepening our understanding of the nature of biological intelligence.
@@ 4 Acknowledgement
We thank M.Murthy et al. for their invaluable help in sharing the Drosophila connectome data. This work was supported by the National Natural Science Foundation of China (72025405, 72088101), the National Social Science Foundation of China (22ZDA102), the Hunan Science and Technology Plan Project (2020TP1013, 2020JJ4673, 2023JJ40685), the Innovation Team Project of Colleges in Guangdong Province (2020KCXTD040), and the Postgraduate Scientific Research Innovation Project of Hunan Province (XJCX2023095). The authors declare that they have no conflict of interest.
@@
Appendix A Related Works
Various models have been developed to mimic or understand aspects of brain function, ranging from the cellular level to the whole brain. In this section, we review the models that aim to simulate the brain’s behavior. The review includes neural dynamic model that describes how action potentials in neurons are initiated and propagated. In addition, we introduce the large scale brain simulating methods. At last, we introduce the network communication models which refers to a computational or theoretical framework designed to understand and simulate how different parts of the brain communicate with each other.
A.1 Neuron Dynamic Model
Neuronal dynamic model is a mathematical framework designed to understand and simulate the electrical activity and dynamic behavior of neurons. These models aim to capture the fundamental mechanisms by which neurons process and transmit electrical signals through their membranes, often by using differential equations and computational techniques. In past years, neuroscientists have aimed to describing the behavior of neurons by dynamic model. One of the most famous works is the Hodgkin-Huxley (HH) model[18]. It includes a set of differential equations that describe the changes in membrane potential and the dynamics of ion channels. The key variables in the model are the membrane potential $V$, and the gating variables $m$, $n$, $h$ that represent the probability of ion channels being open.
The currents are given by:
$I_{Na}$ $=\bar{g}_{Na}m^{3}h(V-E_{Na}),$ (1)
$I_{K}$ $=\bar{g}_{K}n^{4}(V-E_{K}),$ (2)
$I_{L}$ $=\bar{g}_{L}(V-E_{L}),$ (3)
where $\bar{g}_{Na}$, $\bar{g}_{K}$ and $\bar{g}_{L}$ are the maximum conductances for sodium, potassium, and leak channels. $E_{Na}$, $E_{K}$, and $E_{L}$ are the reversal potentials for sodium, potassium, and leak channels.
The gating variables obey first-order kinetics and are described by the following differential equations:
$\frac{dm}{dt}$ $=\alpha_{m}(1-m)-\beta_{m}m,$ (4)
$\frac{dn}{dt}$ $=\alpha_{n}(1-n)-\beta_{n}n,$ (5)
$\frac{dh}{dt}$ $=\alpha_{h}(1-h)-\beta_{h}h,$ (6)
where $m$, $n$, and $h$ are the gating variables representing the activation and inactivation states of the ion channels. $\alpha_{m}$, $\alpha_{n}$, $\alpha_{h}$, $\beta_{m}$, $\beta_{n}$, and $\beta_{h}$ are voltage-dependent rate constants for the gating variables. These rate constants are defined by empirical equations based on experimental data. HH model accurately describes the electrophysiological processes of neuronal activity through mathematical equations, allowing scientists to understand the fundamental mechanisms of a group of neurons’ behavior at the molecular and cellular levels.
Another classic theory is cable theory[19]. Cable theory is a mathematical model used to understand how electrical signals propagate along neurons, especially in their dendrites and axons. This theory applies the principles of electrical circuits to describe how the passive properties of the neuron influence signal transmission. It’s particularly useful in explaining the conduction of graded potentials, such as synaptic potentials and electrotonic (passive) spread of membrane potential changes.
The core equation of cable theory is derived from the cable equation, which is a form of the second-order linear partial differential equation. It can be expressed as:
$\frac{\partial^{2}V}{\partial x^{2}}=\frac{1}{r_{m}}\left(\frac{\partial V}{\partial t}+\frac{V}{r_{c}}\right),$ (7)
where $V$ is the membrane potential, $x$ is the distance along the cable, $t$ is time, $r_{m}$ is the membrane resistance per unit length, $r_{c}$ is the cytoplasmic (axial) resistance per unit length. This equation describes how the voltage $V$ changes along the length of the neuron $x$ and time $t$.
Cable theory has several important applications in neuroscience, including understanding signal decay, neural modeling and synaptic integration. Despite its usefulness, cable theory has limitations. First, real neurons have complex geometries. Cable theory don’t describe well of these complex geometries. This can lead to inaccuracies in predicting real neuronal behavior. Besides, cable theory primarily addresses passive electrical properties and does not account for active properties like action potentials, which are crucial for neural communication.
Apart from that, researchers proposed many variants of neuron dynamic models[20][21]. These models serve to elucidate the physiological activity of neurons and contribute to the description of neuron behavior.
However, these models has some limitations: these equations often contain a multitude of parameters. In instances where measurement techniques fall short of quantifying each parameter individually, estimation becomes a necessity.
Spiking Neural Networks (SNNs)[22] is designed to optimize the parameters in neural dynamic models. Unlike their traditional counterparts, which process information through continuous values, SNNs utilize discrete events known as spikes. These spikes, brief in duration, mimic the communication method of neurons within the brain, facilitating the transmission of information.
One of the simplest and most commonly used models to describe neuron dynamics in SNNs[22] is the Leaky Integrate-and-Fire (LIF) model[23]. The LIF model captures the essential characteristics of spike generation in neurons. The membrane potential $V(t)$ of a neuron at time $t$ is described by the following differential equation:
$\tau_{m}\frac{dV(t)}{dt}=-[V(t)-V_{rest}]+RI(t),$ (8)
where $\tau_{m}$ is the membrane time constant, which determines how fast the membrane potential decays towards the resting potential $V_{rest}$ in the absence of input. $V_{rest}$ is the resting membrane potential. $R$ is the membrane resistance. $I(t)$ is the input current at time $t$. When $V(t)$ reaches a certain threshold $V_{threshold}$, the neuron fires a spike, and $V(t)$ is reset to $V_{rest}$ or another predefined value. After firing, the neuron may enter a refractory period during which it cannot spike, regardless of the input.
SNNs have enabled researchers to simulate networks comprising millions of neurons, tackling a wide array of tasks ranging from image classification and natural language processing to controlling robotic arms[24][25][26][27].
Despite their advances, SNNs face significant challenges. One primary issue is that the SNN’s structure often draws inspiration from functional brain networks rather than accurately replicating the intricate neural connections observed at the microscale. In terms of the connection method of artificial neurons, SNNs usually use a structure similar to that of ANNs. Consequently, these models resort to architectures akin to those found in artificial neural networks, hindering their ability to authentically emulate brain functions. Additionally, SNNs are marked by their low computational efficiency, with a single forward propagation frequently requiring several seconds to complete. These limitations underscore the difficulties in considering SNNs as the optimal approach for brain simulation in their current state.
As for the large scale brain simulation model aims to simulate the whole brain’s behavior, Eugene et al.[28] presents the first large scale brain simulation model based on experimental measures in several mammalian species. The model is founded upon the global thalamocortical anatomy derived from diffusion tensor imaging of the human brain. It incorporates multiple thalamic nuclei and a six-layered cortical microcircuitry, which is grounded in three-dimensional reconstruction of single neurons from the cat visual cortex. It boasts nearly half a billion synapses, incorporating accurate receptor kinetics, short-term plasticity, and dendritic spike-timing-dependent synaptic plasticity (dendritic STDP). The model can demonstrate behavioral patterns that mimic normal brain activity. This model is a comprehensive attempt to simulate the whole neural system, which is a critical component of the brain involved in sensory processing and consciousness.
Eliasmith et al.[29] introduce a more comprehensive brain simulating model, which consists of 2,500,000 SNN neurons. This model receives only sequences of visual images and produces all its responses through a simulated arm. Despite its simplifications of real structure, the model encapsulates numerous elements of neuroanatomy, neurophysiology, and behavioral psychology, as evidenced by its performance across eight distinct tasks.
As for the standardized platform, Zeng et al. [30] presents BrainCog, a cognitive intelligence engine inspired by the brain. This platform, based on SNN, offers critical infrastructure for the creation of brain-inspired artificial intelligence and simulation of the brain. BrainCog encompasses a variety of biological neurons, encoding methods, learning principles, brain regions, and a combination of hardware and software as its core elements.
However, large scale brain simulations require the simulation of millions or even billions of neurons, posing extremely high demands on computational resources. Even the most powerful supercomputers might need weeks or even months to complete a single simulation. Additionally, as the number of neuron connections and synaptic weights increases, data storage and processing requirements also significantly rise. Besides, although large scale brain simulations aim to simulate neurons and their connections, they often fail to capture all the biological details present in the brain. Factors like neuron morphology, molecular-level dynamics, and changes in gene expression are complex biological phenomena that are difficult to fully integrate into the models. Thus, the simulation results might not fully reflect the real behavior of the brain. At last, building a large-scale brain simulation model requires a vast amount of precise biological data, such as the connectivity between neurons, synaptic strengths, types, and concentrations of neurotransmitters. However, acquiring these data is often very challenging, involving complex and expensive experimental techniques, and it’s difficult to measure comprehensively across the entire brain.
As for the collective action of neuron populations, despite the success of neural system mean field models[31] in areas such as modeling epileptic seizures[32], there is still a lack of models that precisely describe the large-scale activity of whole brain.
A.2 Brain Network Communication Model
The network communication model is a conceptual framework which describes how data is transmitted from one node to another over a network. It breaks down the complex process of communication into simpler, more manageable layers or components. The most widely recognized network communication models are the Open Systems Interconnection (OSI) model and the Internet Protocol Suite, commonly known as the TCP/IP model. Nowadays, scientists apply network communication models to neuron research, which are mostly used to understand how brain networks operate[33, 34, 35].
Early studies have indicated that, similar to social and electrical networks[36], brain networks also exhibit properties such as power-law distributions[37] and the rich-club phenomenon[38]. These results indicate that brain networks contain a wealth of network statistical characteristics.
Researchers designed many brain network models to describe the processes of message propagation in brain. One category of brain network communication models is the routing protocol model. The most famous model of this category is shortest path model[39], which assumes that signals between a pair of nodes travel along the shortest path between them. In the shortest path model, the shortest path length from the starting node $s$ to each node $v$, is initialized as $d(s)=0$ and $d(v)=\infty$; then, process each node in increasing order of $d(v)$, for each node, update the $d(u)$ value of its neighbors $u$. The update way can be represented as follows:
$d(u)=\min(d(u),d(v)+w(v,u)),$ (9)
where $w(v,u)$ is the weight of edge $e_{uv}$.
Information is propagated entirely through the shortest paths in these models, which means in terms of propagation cost, signal transmission is optimal. However, the key disadvantage of the model is these methods require each neuron must have knowledge about the entire network structure[40], which doesn’t correspond to the reality: every neuron only has the message of its neighbor neurons.
Another well-known model is navigation model. It suggests that neurons always send information to the neuron that can be reached with the least communication cost[41]. These models achieve the lowest communication cost. A significant achievement is that some algorithms can identify paths with 70%-100% of optimal signaling efficiency in the functional connectomes of different organisms[42]. However, the main disadvantage of navigation model is it supposes that neurons know the distance between their neighbors and a desired target region. Recently their still has a lack of evidences that reveal how this knowledge would be given to neurons in brain networks.
Another well-known category of models is the diffusion model. The diffusion model posits that signals propagate through random walk dynamics. This process of propagation does not require individual neurons to possess any prior knowledge beyond their immediate neighboring neurons. Among the most classic of these is the random walk model[43]. In the random walk model, signals are transmitted to randomly selected neighboring nodes with a probability proportional to the connection weights until they reach the target node. For unbiased random walk, the function of random walk model is as follows.
$P_{ij}=\frac{1}{k_{i}},$ (10)
where $P_{ij}$ represent the probability of node $i$ choosing node $j$.$k_{i}$is the degree of node $i$.
As for biased random walk, for example, if the transition probability depends on the edge weights or certain properties of the nodes, then the transition probability can be adjusted to
$P_{ij}=\frac{w_{ij}}{\sum_{l\in N(i)}w_{il}},$ (11)
where $w_{ij}$ is the weight of node $i$ to node $j$,$N(i)$is the set of node $i$.
The diffusion process does not require individual neurons to have knowledge about the connectivity of connectome beyond their immediate neighbors. However, studies have shown that the efficiency of signal transmission through random walks is relatively low[44]. So it can’t describe the rapid communication process in the brain.
An effective category is threshold model[45]. The threshold model was originally proposed to characterize collective behavior. In the threshold model, a node’s behavior depends on the behavior of its neighboring nodes. Suppose that within a group, nodes can hold one of two states regarding a certain behavior: an active state and a quiescent state. For a node, it will only transition from the quiescent state to the active state if the stimulus given by its neighboring nodes in the active state surpasses a certain threshold; otherwise, it remains in the quiescent state. The condition for node $i$ to adopt the behavior at time $t$+1 can be formulated as follows.
$a_{i}(t+1)=1\quad\text{if}\quad\frac{1}{k_{i}}\sum_{j\in N_{i}}a_{j}(t)\geq\theta_{i},$ (12)
where $N_{i}$ is the set of neighbors of node $i$, $a_{i}(t)$ is the adoption state of node $i$ at time $t$. $\theta_{i}$ is the threshold of node $i$. $k_{i}$ is the degree of node $i$. The threshold model has stronger biological interpretability because it approximates the mechanism of neuronal firing. Consequently, the threshold model has been extensively applied in artificial neural networks. However, how to set the propagation thresholds for different neurons, and how to design the dynamics model between neurons remains a challenge.
Appendix B Data and Methods
B.1 Drosophila Connectome Data Set
@@
We employ the recently released largest Drosophila brain connectome[17], offering a thorough mapping of the entire neural network within a Drosophila female brain. This data set encompasses over 120,000 neurons and 30,000,000 synapses, including detailed information on synaptic connections between neurons. Notably, it provides high-precision 3D coordinates, accurate to the nanometer level, for both cell bodies and synapses. Additionally, this data set offers extensive labeling for various types of neurons. For a more in-depth understanding, please refer to [17][46] for detailed information. In this section, we use nodes to represent neurons and connections to represent the synapses between neurons.
@@
Table 3 presents the basic properties of the data set. Note that we combine multiple edges between two nodes to calculate network statistic properties. The average degree before and after combining multiple edges is 501.6 and 37.01, respectively, indicating it is prevalent that the pairs of neurons are connected by multiple synapses.
@@
Network Diameter Assortativity Coefficient
$11$ $0.0008817$
B.2 Model

### v4

@@ 2.4 Distances between Input and Activation Areas
To further investigate whether this separation phenomenon is caused by greater spatial distance, we calculated the average network distance and spatial distance between different areas of the brain network. The results are presented in Tab. 1.
The analysis of Tab. 1 showcases that spatial distances among neurons do not straightforwardly translate to ease of activation. While visual input neurons across all intrinsic areas exhibit similar average spatial distances, quasi-real activation patterns manifest predominantly in specific neural types (optic and visual_projection). Conversely, the average network distances between input neurons and target neurons is notably smaller compared to irrelevant neurons. Specifically, visual input neurons exhibit network distances to optic and visual_projection neurons as 5.7018 and 5.7432, respectively, while distances to neurons irrelevant to olfactory (ALLN, ALIN, ALON, and ALPN) are measured at 6.6442, 6.2350, 6.1506, and 6.5380, respectively. This demonstrates that neurons corresponding to visual inputs maintain a generally closer network distance relative to unrelated neurons. Furthermore, neurons associated with olfactory inputs maintain network distances approximately half that of unrelated neurons. These observations suggest that the network distance between input neurons and their specific activation areas is more significant than spatial distance, underscoring the importance of network structure over spatial structure in defining the quasi-real activation patterns of brain networks.
Table 1: Spatial and network distance between input areas and functional areas.
Input type Distance type optic visual_projection ALLN ALON ALPN ALIN
Visual input Spatial distance (nm) $365012$ $354725$ $351919$ $346925$ $348517$ $348901$
Network distance (edge) $\mathbf{5.7018}$ $\mathbf{5.7432}$ $6.6442$ $6.1506$ $6.5380$ $6.2350$
Olfactory input Spatial distance (nm) $302690$ $251144$ $110827$ $123179$ $115089$ $173345$
Network distance (edge) $5.7967$ $5.1706$ $\mathbf{3.4458}$ $\mathbf{3.6312}$ $\mathbf{3.6820}$ $\mathbf{3.3845}$
@@ 3 Discussion
Our study, which utilized a large-scale network communication model derived from the Drosophila connectome, highlights the crucial importance of neural network architecture in generating brain-like activation patterns.
We first analyzed the structure of the Drosophila brain network, including the basic statistical properties and the symmetry of the network. An interesting phenomenon observed is the correlation between neuronal degree and geometric size: Fig. 1(f) demonstrates that the number of synapses formed by a neuron (represented by degree) increases logarithmically, rather than linearly, with respect to its surface area, volume, or length. These characteristics of the brain can be explained by the Weber-Fechner’s law: for any sensory pattern, the perceived intensity is a logarithmic function of the physical intensity. Specifically, the size, length, and area changes can only cause a logarithmic change in the number of synapses the neuron has. These results suggest that although synaptic plasticity lead to a continuous increase in the number of synapses between neurons and other cells, this logarithmic relationship makes it difficult for the number of synapses formed by a neuron to increase indefinitely. This may account for the overall power-law distribution observed in Fig. 1(f).
To examine the impact of neuronal activation mechanisms on the generation of realistic network patterns, we simulated three activation mechanisms for specific inputs. Fig. 2 suggests that, the real network structure is more critical than activation mechanisms in generating quasi-real activation patterns. The brain’s unique network structure plays a crucial role in this realistic activation mechanism, aligning with studies on complex systems where the organization and interaction structure among individuals, rather than individual capabilities, determine emergent phenomena[14].
To elucidate the influence of network structure on function, we analyzed the topology of the network (Fig. 3 and Tab. 1). These results collectively suggest that network architecture, rather than spatial structure or activation mechanisms, is the primary determinant of the Drosophila connectome’s ability to generate quasi-real activation patterns, owing to its unique connectivity profile.
To test the impact of network perturbation ratio on its functionality, we calculated activation patterns with different reconnect rates (Appendix S5). These results indicate that the formation of realistic activation patterns is highly sensitive to changes in network structure, with even minor perturbations capable of disrupting this network configuration. This indirectly substantiates that the authentic network structure is crucial for the formation of genuine activation patterns. Additionally, it highlights that complex functions (such as vision) are more susceptible to network disturbances. Conversely, Fig. S4 shows that the network retains its original performance under a region-based attack. One explanation for these findings is that random rewiring thoroughly alters the network topology, potentially disrupting original functional modules and pathways, which leads to disordered information transmission paths and significantly impacts activation patterns. Additionally, because rewiring is global and random, it induces unpredictable perturbations that undermine redundancy and fault tolerance mechanisms, rendering the network more sensitive to such global disturbances. In contrast, the reason the network’s dynamic properties remain relatively stable after the deletion of nodes within a specific region is that this perturbation is localized and has a minor impact on the global topology. Moreover, random rewiring may affect critical nodes or key pathways within the network, leading to significant functional impairment of the overall network. However, regional node deletion, if involving non-critical nodes and edges, may not significantly impact the network’s functionality, allowing it to maintain normal activation patterns.
To evaluate the impact of the observed network structural symmetry (Fig. 1(a), (c), (d), and (e)), we input visual signals separately to the left and right hemispheres and observed the effects. The findings in Appendix D reflect real-world phenomena: Although the left and right hemispheres possess certain functional specializations, they also exhibit synergistic effects and typically collaborate to fulfill specific functions [15]. For example, during complex cognitive tasks, both hemispheres coordinate their efforts to achieve task completion. The experimental results indicate that similar synergistic effects emerge even when only the neurons on one side of the brain are stimulated, indirectly validating the model’s effectiveness. Moreover, the network maintains similar synergistic effects even in the face of large-scale localized attacks, indirectly demonstrating the network’s robust resilience. Finally, by analyzing the connections between different regions, it is found that the cause of these synergistic effects is not the connections between visual input and contralateral visual processing areas, but rather the extensive connections within the visual processing regions themselves. This provides a new perspective for understanding the brain’s synergistic effects.
@@
These findings has significant implications for artificial neural network research, suggesting that future AI designs may need to more closely mimic real brain network structures. Future research directions may include optimizing propagation rules and automatically adjusting link weights to achieve complete simulation of brain behavior. These advancements will pave the way for achieving genuine artificial intelligence while deepening our understanding of the nature of biological intelligence.
@@ 4 Acknowledgement
We thank M.Murthy et al. for their invaluable help in sharing the Drosophila connectome data. This work was supported by the National Natural Science Foundation of China (72025405, 72088101, 72301285, 72001211), the National Social Science Foundation of China (22ZDA102), and the Natural Science Foundation of Hunan Province (2023JJ40685). The authors declare that they have no conflict of interest.
@@
Appendix A Data and Methods
A.1 Drosophila Connectome Data Set
@@
We utilize the recently released largest Drosophila brain connectome[12], which provides a comprehensive mapping of the entire neural network within a female Drosophila brain. This data set includes over 120,000 neurons and 30,000,000 synapses, featuring detailed synaptic connections and high-precision 3D coordinates (accurate to the nanometer level) for both cell bodies and synapses. Additionally, it offers extensive labeling for various neuron types. For more detailed information, refer to [12, 16].
@@
In this section, neurons are represented as nodes and synapses as connections between them. Tab. 2 presents the basic properties of the data set. We combine multiple edges between two nodes to calculate network statistics. The average degree before and after combining multiple edges is 501.6 and 37.01, respectively, indicating the prevalence of multiple synapses between pairs of neurons.
@@
Network Diameter Average Shortest Path Assortativity Coefficient
$11$ $3.9061$ $-0.05868$
A.2 Model